Damarion4:

escribing Effects of a Change in Slope and y-Intercept The yearly cost in dollars, y, at a video game arcade based on total game tokens purchased, x, is y = y equals StartFraction 1 Over 10 EndFraction x plus 60.x 60 for a member and y = y equals StartFraction 1 Over 5 EndFraction x. x for a nonmember. Explain how the graph of a nonmember’s yearly cost will differ from the graph of a member’s yearly cost.

aarondkdkjd:

cant really understand this

aarondkdkjd:

a bunch of start fractiions and over 10

aarondkdkjd:

can you rewrite this

TheSmartOne:

For a member: \[y = \frac{ 1 }{ 10}x + 60\] For a nonmember \[y = \frac{1}{5}x \] Would you like to help explain it now for the asker, @aarondkdkjd